# Two-dimensional inference of divertor plasma characteristics: advancements to a multi-instrument Bayesian analysis system

### v1

@@ 3.1 A Field-Aligned Inference Mesh
As with [23], we seek a mesh-based inference. Much of the physics in divertor plasmas is tied to magnetic flux surfaces and consequently is spatially-dependent. Bayesian inference permits a priori information to our inference in the form of the prior probability of equation (1). Consequently, to enable an efficient manner for including prior information on anisotropic plasma properties, we align our inference mesh to surfaces of constant poloidal magnetic flux.
@@ 3.2 The Posterior Distribution
The logarithm of the posterior distribution of equation (1), $\mathcal{L}_{\mathrm{Posterior}}$, is the sum of the log-likelihood, $\mathcal{L}_{\mathrm{Likelihood}}$, log-prior, $\mathcal{L}_{\mathrm{Prior}}$ and a normalisation constant, $c$, of which is omitted due to its lack of influence on either the MAP estimate or posterior distribution width (uncertainty). Both the log-likelihood and log-prior are themselves the sum of numerous log-probability distributions, respectively pertaining to different diagnostics, $i$, and different physics-informed priors, $j$. The total log-posterior probability is therefore,
@@ 3.3 Likelihoods
For each diagnostic, a multivariate Gaussian likelihood distribution (uncorrelated between diagnostics) was assumed such that each diagnostics’ log-likelihood (minus an omitted normalisation constant) was given by
@@
$\mathbf{d}_{i}$ specifies a vector of measurements for the $i^{th}$ diagnostic. $f_{i}\left(\mathbf{\theta}\right)$ represents the forward model that predicts a diagnostic response from a parameter state (represented by the model parameters, $\theta$). $\mathbf{\Sigma}$ is the covariance matrix which contains information on uncertainties.
@@
  • camera images  [^3] Modelled after MAST-U’s Multi-Wavelength-Imaging (MWI) diagnostic [41, 24] and TCV’s MANTIS diagnostic [20] of the brightness of
@@
Thomson scattering measurements of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$, along with uncertainties, are routinely available at MAST-U [42]. However, data points may be unavailable or not perfectly aligned to specific flux tubes, which we reflect in our chosen positions shown in figure 1. We assume the availability of Langmuir probe $j_{\mathrm{sat}}$ measurements along MAST-U’s tile 5 (figure 1). Assuming similar ion and electron temperatures in the target vicinity, we model
@@
where $e$ is the elementary charge and $m_{i}$ is the ion mass (taken exclusively as deuterium ions). We assume the availability of brightness images (calibrated spatially and in absolute brightness value) for specific wavelength windows associated with each line. To model each line’s brightness image, the line’s emissivity was modelled at each mesh vertex and a geometry matrix was used to transform the poloidal cross section emissivity into camera brightness.
@@ 3.3.1 Emission models
where $i$ is a contributing process to the $m\to n$ emission. $n_{i,0}$ and $n_{i,1}$ are the number densities of the two species involved in the process. The $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ dependent photon-emissivity coefficient, PEC, of the process is generated from generalised collisional-radiative models (provided by ADAS [22] for atomic interactions).
@@ 3.3.2 Inclusion of molecular effects in the hydrogen Balmer emission model
Hydrogenic Balmer line emission is typically modelled through two predominant atomic processes: electron-impact recombination (EIR) and electron-impact excitation (EIE). However, recent work [9, 26] has shown that hydrogen molecules can play an important role in both monitored hydrogen emission and plasma dynamics through molecular activated recombination (MAR) and molecular activated dissociation (MAD). MAR and MAD create excited hydrogen atoms and so PMI must be considered as an additional process in  (5).
A comprehensive model of PMI contributions to Balmer lines requires knowledge of $D_{2}$ plasma chemistry. To avoid the need to introduce the various parameters specifying this plasma chemistry, we parameterise molecular contributions to Balmer line emission via a single parameter, $Q_{\mathrm{mol}}$, in the appendix A.1. $Q_{\mathrm{mol}}$ is defined as the ratio of $D_{\alpha}$’s molecular emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{mol}}$, to atomic emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{atm}}$,
@@
As outlined in the appendix A.1, an effective molecular PEC, $\mathrm{PEC}{}^{n\to 2}_{D{2,\mathrm{eff}}{}}$, was generated to conglomerate the various molecular contributions to a line’s transition. The emissivity of Balmer lines $n\to 2$ were thus given by,
$\eqalign{\varepsilon_{n\to 2}=\underbrace{n_{e}^{2}\mathrm{PEC}{}^{n\to 2}_{eir}+n_{0}n_{e}\mathrm{PEC}{}^{n\to 2}_{\mathrm{eie}}}_{\mathrm{Atomic\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{atm}}}+\cr\ \underbrace{Q_{D_{\alpha}}^{\mathrm{mol}}\left(n_{e}^{2}\mathrm{PEC}{}^{3\to 2}_{\mathrm{eir}}+n_{0}n_{e}\mathrm{PEC}{}^{3\to 2}_{\mathrm{eie}}\right)\frac{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\to 2}}{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{3\to 2}}}_{\mathrm{Molecular\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{mol}}}.}$ (7)
@@ 3.3.3 Helium emission model
MAST-U’s MWI and TCV’s MANTIS cameras also routinely capture Helium-I emission lines which has been shown [40] to provide information on $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$. To avoid inaccuracies in Helium camera forward models, we explicitly include parameters for the densities of neutral and singly charged helium ($n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ respectively) at each mesh vertex (at a cost of 2V additional free parameters) in our inference. This allowed both EIE and EIR contributions to be modelled via equation (5). Following the work of [43], 668, 728 and 502 nm singlet He I lines were chosen due to their insensitivity to transport of metastable states. Additional details and potential complexities are outlined in the appendix A.2.
@@ 3.4 Priors
We refer to each physics-based constraint considered in the inference as a prior. Each prior is imposed as a contribution, $\mathcal{L}_{j}\left(\theta\right)$, to the overall log-prior of equation (2) with details of the probability function used provided in the appendix C.1. Each prior is evaluated at all mesh vertices relevant to that prior.
@@ 3.4.1 Spatially-Independent and -Dependent Priors
We consider ‘spatially-independent priors’ to be those that, when being evaluated at a certain mesh vertex, require no information about parameter values at a different mesh vertex. Consequently, these can be included in both the cell-based and mesh-based inference paradigms. They include priors that consider the interplay between parameters at a single mesh vertex such as an upper bound of the static electron pressure as well as bounds on parameter values. All spatially-independent priors used in this work are detailed in the appendix C.3.
‘Spatially-dependent priors’, when being evaluated at a certain mesh vertex, require information of parameter values at other mesh vertices and so cannot be used in the cell-based inference paradigm. An example of such a prior is the expected monotonic reduction (toward the target) in the heat flux parallel to a surface of constant poloidal magnetic flux. All spatially-dependent priors used in this work are detailed in the appendix C.4.
@@ 3.5 Inference Parameters
In the mesh based inference, the model parameters,
@@
$F=\left\{T_{e},n_{e},n_{0},Q_{\mathrm{mol.}},n_{He^{0}},n_{He^{1+}}\right\}.$ (9)
$T_{e},n_{e},n_{0}$ are of direct interest for divertor physics. These, as well as $Q_{\mathrm{mol}}$, can be used to derive further quantities of interest to divertor physics including: static electron pressure; heat flux parallel to flux tubes, divertor ionisation sources; ion sinks (volumetric recombination due to electron-ion recombination and molecular activated recombination); hydrogenic power losses; and more [44]. $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ parameters can be identified as nuisance parameters which are necessary to evaluate forward models but not of direct interest in this work.
@@ 4 Synthetic Data Setup
(b) Maximum a posteriori (MAP) estimate inference of scenarios displayed in Figure 2(a) using posterior i. MAP was found using the generational algorithm outlined in appendix A of [23].
The new integrated data analysis methodology developed in this work was applied to synthetic data. The data were generated from synthetic diagnostics applied to two SOLPS-ITER simulations: a ‘detached case’, shown in figure 2aI, where the plasma was in a detached state in which PMI was expected to play a significant role; and an ‘attached case’, shown in figure 2bII, where the plasma was in an attached state in which PMI was not expected to play a significant role. These simulations had an input power of 2.5 MW and were obtained from [45] using the ‘Super-X low $\alpha$’ geometry  [^4] This geometry is observed more easily by the synthetic imaging diagnostic as the Super-X has no curvature in the separatrix near the target, see figure 2 of [45]. A trace amount of helium was added to get realistic relative 2D spatial distributions of $n_{{\mathrm{He}}^{{0}}}$ and its ratio to $n_{{\mathrm{He}}^{{1+}}}$, although the absolute numbers for $n_{{\mathrm{He}}^{{0}}}$ may not be characteristic of that obtained in the tokamak  [^5] However, the absolute magnitude of $n_{{\mathrm{He}}^{{0}}}$ is not expected to impact our inference..
For each case, the SOLPS-ITER simulation was taken as the ‘true’ plasma state from which the relevant fields were extracted. To compute $Q_{\mathrm{mol}}$, $\mathrm{PEC}_{\mathrm{D{2,eff}}}$ was used to post-process the $\mathrm{D}_{\mathrm{2}}$ density into the densities for the various molecular ions. It should be noted that this leads to significantly higher $\mathrm{D}_{\mathrm{2}}^{+}/\mathrm{D}_{\mathrm{2}}$ ratios than reported by SOLPS-ITER directly, which is known to under-predict the $\mathrm{D}_{\mathrm{2}}^{+}/\mathrm{D}_{\mathrm{2}}$ ratio due to inaccuracies in the molecular charge exchange rate (see [46, 10, 27] for more information).
@@ 5 Results
We have applied the D-MIBAS framework to the two synthetic data sets outlined in section 4 (with both including PMI in the measured Balmer-line brightness). Four different posterior distributions have been constructed under the framework to ascertain the influence of the advanced emission model, the use of spatially-dependent priors and the inclusion of additional diagnostics. These posterior distributions are:
@@ 5.1 Synthetic diagnostic test results of the full D-MIBAS analysis
Figure 3: Comparison of true and inferred (MAP) parameter values using posterior i for case detached case. The 95% highest density interval (HDI) for the inferred parameters are displayed. The percentage of true data points lying within the 95% HDI are 95%, 88%, 63%, 91%, 95%, 63% for the $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, $n_{\mathrm{H}}$, $Q_{\mathrm{mol}}$, $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ respectively.
@@
The inferred uncertainties in the parameters (their 95% HDI) were found to be largely in agreement with the error between the true parameter values and the inferred MAP estimates. This is demonstrated in figure 3. The uncertainty was found to be substantially greater for $n_{\mathrm{H}}$ and $Q_{\mathrm{mol}}$ parameters, (as well as the nuisance parameters $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$) compared to $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ parameters. $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ parameters with large uncertainty were found to be in regions out of view of the synthetic camera diagnostic (upstream) and in regions of low Balmer and He I emission (deep into both the private flux and common flux regions).
@@ 5.2 Performance of the different posterior distributions
Inference with each posterior distribution was able to accurately recreate the synthetic data, as shown in figure 4. However, the inferred parameters (MAP estimate) of each posterior distribution were considerably different. The resultant MdAPE scores of inference with each of the posterior distributions outlined in section 5 are shown in figure 5. All parameters of interest were inferred most accurately by the most complete version of D-MIBAS (posterior i). We will now discuss each reduced analysis case (posteriors ii-iv) in more detail.

### v2

@@ 3.1 A Field-Aligned Inference Mesh
As with [Bowman_2020], we seek a mesh-based inference. Much of the physics in divertor plasmas is tied to magnetic flux surfaces and consequently is spatially-dependent. Bayesian inference permits a priori information to our inference in the form of the prior probability of equation (1). Consequently, to enable an efficient manner for including prior information on anisotropic plasma properties, we align our inference mesh to surfaces of constant poloidal magnetic flux.
@@ 3.2 The Posterior Distribution
The logarithm of the posterior distribution of equation (1), $\mathcal{L}_{\mathrm{Posterior}}$, is the sum of the log-likelihood, $\mathcal{L}_{\mathrm{Likelihood}}$, log-prior, $\mathcal{L}_{\mathrm{Prior}}$ and a normalisation constant, $c$, which is neglected due to its lack of influence on either the MAP estimate or posterior distribution width (uncertainty). Both the log-likelihood and log-prior are themselves the sum of numerous log-probability distributions, respectively pertaining to different diagnostics, $i$, and different physics-informed priors, $j$. The total log-posterior probability is therefore,
@@ 3.3 Likelihoods
For each diagnostic, a multivariate Gaussian likelihood distribution (uncorrelated between diagnostics) was assumed such that each diagnostic log-likelihood (minus an omitted normalisation constant) was given by
@@
$\mathbf{d}_{i}$ specifies a vector of measurements for the $i^{th}$ diagnostic. $f_{i}\left(\mathbf{\theta}\right)$ represents the forward model that predicts a diagnostic response from a plasma state (represented by the model parameters, $\theta$). $\mathbf{\Sigma}$ is the covariance matrix, which contains information on uncertainties.
@@
  • camera images  [^4] Modelled after MAST-U’s Multi-Wavelength-Imaging (MWI) diagnostic [Feng2017, Wijkamp2023] and TCV’s MANTIS diagnostic [Perek2019submitted] of the brightness of
@@
Thomson scattering measurements of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$, along with uncertainties, are routinely available at MAST-U [Hawke_2013]. However, data points may be unavailable or not perfectly aligned to specific flux tubes, which we reflect in our chosen positions shown in Figure 1. We assume the availability of Langmuir probe $j_{\mathrm{sat}}$ measurements along MAST-U’s tile 5 (Figure 1). Assuming similar ion and electron temperatures in the target vicinity, we model
@@
where $e$ is the elementary charge and $m_{i}$ is the ion mass (taken exclusively as deuterium ions). We assume the availability of brightness images (calibrated spatially and in absolute brightness value) for specific wavelength windows associated with each line. To model each line’s brightness image, the line’s emissivity was modelled at each mesh vertex and a geometry matrix was used to transform the poloidal cross-section emissivity into camera brightness.
@@ 3.3.1 Emission models
where $i$ is a contributing process to the $m\to n$ emission. $n_{i,0}$ and $n_{i,1}$ are the number densities of the two species involved in the process. The $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ dependent photon-emissivity coefficient, PEC, of the process is generated from generalised collisional-radiative models (provided by ADAS [adas] for atomic interactions).
@@ 3.3.2 Inclusion of molecular effects in the hydrogen Balmer emission model
Hydrogenic Balmer line emission is typically modelled through two predominant atomic processes: electron-impact recombination (EIR) and electron-impact excitation (EIE). However, recent work [Verhaegh_2022, Karhunen2023] has shown that hydrogen molecules can play an important role in both monitored hydrogen emission and plasma dynamics through molecular activated recombination (MAR) and molecular activated dissociation (MAD). MAR and MAD create excited hydrogen atoms and so PMI must be considered as an additional process in equation (5).
A comprehensive model of PMI contributions to Balmer lines requires knowledge of $D_{2}$ plasma chemistry. To avoid the need to introduce the various parameters specifying this plasma chemistry, we parameterise molecular contributions to Balmer line emission via a single parameter, $Q_{\mathrm{mol}}$, in Appendix A.1. $Q_{\mathrm{mol}}$ is defined as the ratio of $D_{\alpha}$’s molecular emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{mol}}$, to atomic emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{atm}}$,
@@
As outlined in Appendix A.1, an effective molecular PEC, $\mathrm{PEC}{}^{n\to 2}_{D{2,\mathrm{eff}}{}}$, was generated to conglomerate the various molecular contributions to a line’s transition. The emissivity of Balmer lines $n\to 2$ was thus given by,
$\eqalign{\varepsilon_{n\to 2}=\underbrace{n_{e}^{2}\mathrm{PEC}{}^{n\to 2}_{eir}+n_{H}n_{e}\mathrm{PEC}{}^{n\to 2}_{\mathrm{eie}}}_{\mathrm{Atomic\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{atm}}}+\cr\ \underbrace{Q_{D_{\alpha}}^{\mathrm{mol}}\left(n_{e}^{2}\mathrm{PEC}{}^{3\to 2}_{\mathrm{eir}}+n_{H}n_{e}\mathrm{PEC}{}^{3\to 2}_{\mathrm{eie}}\right)\frac{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\to 2}}{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{3\to 2}}}_{\mathrm{Molecular\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{mol}}}.}$ (7)
@@ 3.3.3 Helium emission model
MAST-U’s MWI and TCV’s MANTIS cameras also routinely capture Helium-I emission lines which have been shown [Linehan2023] to provide information on $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$. To avoid inaccuracies in Helium camera forward models, we explicitly include parameters for the densities of neutral and singly charged helium ($n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ respectively) at each mesh vertex (at a cost of 2V additional free parameters) in our inference. This allowed both EIE and EIR contributions to be modelled via equation (5). Following the work of [Biggelaar2022], 668, 728 and 502 nm singlet He I lines were chosen due to their insensitivity to the transport of metastable states. Additional details and potential complexities are outlined in Appendix A.2.
@@ 3.4 Priors
We refer to each physics-based constraint considered in the inference as a prior. Each prior is imposed as a contribution, $\mathcal{L}_{j}\left(\theta\right)$, to the overall log-prior of equation (2) with details of the probability function used provided in Appendix C.1. Each prior is evaluated at all mesh vertices relevant to that prior.
@@ 3.4.1 Spatially-Independent and -Dependent Priors
We consider ‘spatially-independent priors’ to be those that, when being evaluated at a certain mesh vertex, require no information about parameter values at a different mesh vertex. Consequently, these can be included in both the cell-based and mesh-based inference paradigms. They include priors that consider the interplay between parameters at a single mesh vertex such as an upper bound of the static electron pressure as well as bounds on parameter values. All spatially-independent priors used in this work are detailed in Table 1 of Appendix C.2.
‘Spatially-dependent priors’, when being evaluated at a certain mesh vertex, require information of parameter values at other mesh vertices and so cannot be used in the cell-based inference paradigm. An example of such a prior is the expected monotonic reduction (toward the target) in the heat flux parallel to a surface of constant poloidal magnetic flux. All spatially-dependent priors used in this work are detailed in Table 2 of Appendix C.2.
@@ 3.5 Inference Parameters
In the mesh-based inference, the model parameters,
@@
$F=\left\{T_{e},n_{e},n_{H},Q_{\mathrm{mol.}},n_{He^{0}},n_{He^{1+}}\right\}.$ (9)
$T_{\mathrm{e}}$, $n_{\mathrm{e}}$ and $n_{\mathrm{H}}$ are of direct interest for divertor physics. These, as well as $Q_{\mathrm{mol}}$, can be used to derive further quantities of interest to divertor physics including: static electron pressure; heat flux parallel to flux tubes; divertor ionisation sources; ion sinks (volumetric recombination due to electron-ion recombination and molecular activated recombination); hydrogenic power losses; and more [Verhaegh2021a]. $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ parameters can be identified as nuisance parameters which are necessary to evaluate forward models but not of direct interest in this work.
@@ 4 Synthetic Data Setup
(b) Maximum a posteriori (MAP) estimate inference of scenarios displayed in Figure 2(a) using posterior i. MAP was found using the generational algorithm outlined in appendix A of [Bowman_2020].
The new integrated data analysis methodology developed in this work was applied to synthetic data. The data were generated from synthetic diagnostics applied to two SOLPS-ITER simulations: a ‘detached case’, shown in Figure 2aI, where the plasma was in a detached state in which PMI was expected to play a significant role; and an ‘attached case’, shown in Figure 2bII, where the plasma was in an attached state in which PMI was not expected to play a significant role. These simulations had an input power of 2.5 MW and were obtained from [Fil_2022] using the ‘Super-X low $\alpha$’ geometry  [^5] This geometry is observed more easily by the synthetic imaging diagnostic as the Super-X has no curvature in the separatrix near the target, see Figure 2 of [Fil_2022].. A trace amount of helium was added to get realistic relative 2D spatial distributions of $n_{{\mathrm{He}}^{{0}}}$ and its ratio to $n_{{\mathrm{He}}^{{1+}}}$, although the absolute numbers for $n_{{\mathrm{He}}^{{0}}}$ may not be characteristic of that obtained in the tokamak  [^6] However, the absolute magnitude of $n_{{\mathrm{He}}^{{0}}}$ is not expected to impact our inference..
For each case, the SOLPS-ITER simulation was taken as the ‘true’ plasma state from which the relevant fields were extracted. To compute $Q_{\mathrm{mol}}$, $\mathrm{PEC}_{\mathrm{D{2,eff}}}$ was used to post-process the $\mathrm{D}_{\mathrm{2}}$ density into the densities for the various molecular ions. It should be noted that this leads to significantly higher $\mathrm{D}_{\mathrm{2}}^{+}/\mathrm{D}_{\mathrm{2}}$ ratios than reported by SOLPS-ITER directly, which is known to under-predict the $\mathrm{D}_{\mathrm{2}}^{+}/\mathrm{D}_{\mathrm{2}}$ ratio due to inaccuracies in the molecular charge exchange rate (see [Verhaegh2021, Verhaegh2023c, Verhaegh2023a]).
@@ 5 Results
We have applied the D-MIBAS framework to the two synthetic data sets outlined in Section 4 (with both including PMI in the measured Balmer-line brightness). Four different posterior distributions have been constructed under the framework to ascertain the influence of the advanced emission model, the use of spatially-dependent priors, and the inclusion of additional diagnostics. These posterior distributions are:
@@ 5.1 Synthetic diagnostic test results of the full D-MIBAS analysis
Figure 3: Comparison of true and inferred (MAP) parameter values using posterior i for the detached case. The 95% highest density interval (HDI) for the inferred parameters are displayed. The percentage of true data points lying within the 95% HDI are 95%, 88%, 63%, 91%, 95%, 63% for the $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, $n_{\mathrm{H}}$, $Q_{\mathrm{mol}}$, $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ respectively.
@@
The inferred uncertainties in the parameters (their 95% HDI) were found to be largely in agreement with the error between the true parameter values and the inferred MAP estimates. This is demonstrated in Figure 3. The uncertainty was found to be substantially greater for $n_{\mathrm{H}}$ and $Q_{\mathrm{mol}}$ parameters, (as well as the nuisance parameters $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$) compared to $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ parameters. $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ parameters with large uncertainty were found to be in regions out of view of the synthetic camera diagnostic (upstream) and in regions of low Balmer and He I emission (deep into both the private flux and common flux regions). Figure 9 of Appendix B.2 further demonstrates the inferred uncertainty was able to reflect the error in the inferred parameters.
@@ 5.2 Performance of the different posterior distributions
Inference with each posterior distribution was able to accurately recreate the synthetic data, as highlighted in Figure 4. However, the inferred parameters (MAP estimate) of each posterior distribution were considerably different. The resultant MdAPE scores of inference with each of the posterior distributions outlined in Section 5 are shown in Figure 5. All parameters of interest were inferred most accurately by the most complete version of D-MIBAS (posterior i). We will now discuss each reduced analysis case (posteriors ii-iv) in more detail.